OpenStudy (anonymous):

how does 3^5x/2 (times) 3^12x-3 turn into 3^29/2x-3?

OpenStudy (anonymous):

\[3^{\frac{5x}{2}}\times 3^{{12x-3}}\]?

OpenStudy (anonymous):

if so, your job is only to add the exponents: \[\frac{5x}{2}+12x-3=\frac{5x+24x-6}{2}=\frac{29x-6}{2}=\frac{29x}{2}-3\]\]

OpenStudy (anonymous):

so the answer is x=3/10 which i totally get, but my professor went from 3^5x...blah blah wa i wrote earlier (just like u wrote it) to the 29/2x-3 which is wa i dont get?

OpenStudy (anonymous):

i knw i knw sounds confusing